Calculate the number of joules given off when 32.0 grams of steam cools from 110.0 °C to ice at -40.0 °C.


1
Expert's answer
2021-06-03T09:16:54-0400

First the steam should cool to 100°C, next it should turn into water and cool to 0°C, then the water should turn into ice and the ice should cool to -40.0°C. Let us write the amount of energy corresponding to each stage of process.

1) "Q_1 = c_s m_s \Delta T_s = 2078\cdot 0.032\cdot 10°\mathrm{C} = 665\,\mathrm{J}"

2) "Q_2 = r m_s = 2.26\cdot10^6\cdot0.032= 72320\,\mathrm{J}"

3) "Q_3 = c_w m_w \Delta T_w = 4200\cdot0.032\cdot100°\mathrm{C} = 13440\,\mathrm{J}"

4) "Q_4 = \lambda m_w = 330000\cdot0.032 = 10560\,\mathrm{J}"

5) "Q_5 = c_im_i\Delta T_i = 2110\cdot0.032\cdot40°\mathrm{C} = 2701\,\mathrm{J}."

The total amount of energy will be "665+72320+13440+10560+2701 =99686\,\mathrm{J}\approx 10^5\,\mathrm{J}."